# Think, Speak, Decide: Language-Augmented Multi-Agent Policy Learning in Economic Environments

### v1

@@ 1 Introduction
To address this challenge, we propose LAMP (Language‑Augmented Multi‑Agent Policy Learning), which integrates LLM‑driven reasoning and reflection over both numerical observations and textual signals to support optimal decision‑making. LAMP follows a unified Think–Speak–Decide pipeline: (1) Think: Agents receive environment observations and generate both short‑term shock analysis and long‑term trend reasoning via an LLM. High‑reward reasoning trajectories are stored in an experience pool for retrieval in similar contexts. The long‑term reasoning is also passed to the Speak module to inform message generation. (2) Speak: Guided by the Think module, each agent formulates multiple candidate public messages. A lightweight attention‑based scorer selects one for broadcast. Other agents parse the message via the LLM, updating their beliefs, trust, and reflective states. These updated reflections are then passed to the Decide module. (3) Decide: The policy network integrates numerical observations, Think’s reasoning outputs, and Speak’s reflections into the RL policy. Under centralized training with a shared critic, agents learn strategies capable of processing reasoning and reflection signals to produce robust, language‑aware economic decisions. We evaluate LAMP in TaxAI: it outperforms MARL and LLM-only baselines with up to 63% higher returns and 55% better shock robustness. Its reasoning traces explain language-guided choices, aiding insight and policy.
@@
5:   Generate news: $\mathcal{R}^{type}_{t}\leftarrow\textsc{Think}(\cdot)_{type}$
@@
11:     $\mathcal{L}_{\text{reason}}(\mathcal{R}^{type}_{t},O_{t}^{h,i},\mathcal{H}_{k,t}^{i})$
@@
13:      Generate statement: $v_{t}^{i}\leftarrow\textsc{Speak}(O_{t}^{h,i},\mathcal{R}_{t},\text{)}$
@@ 3.1 Problem Formulation
We then model the economic decision‑making problem as a partially observable Markov game $\mathcal{M}=\bigl\langle N,\,S,\,O,\,A,\,P,\,R,\,\gamma\bigr\rangle,$ where $N=\{1,\dots,N_{h}\}$, $\gamma\in[0,1)$, and $P$ is the transition kernel induced by $A=A^{g}\times A^{h,1}\times\cdots\times A^{h,N_{h}}.$ At each step, the government’s action is $A_{t}^{g}=\{\tau_{t},\ \xi_{t},\ \tau_{a,t},\ \xi_{a,t},\ r_{t}^{G}\},$ where $\tau_{t}$ and $\xi_{t}$ parameterize the marginal income-tax schedule, $\tau_{a,t}$ and $\xi_{a,t}$ analogously parameterize the marginal asset-tax schedule, and $r_{t}^{G}$ denotes the expenditure-to-output ratio. Each household $i$ selects a savings rate and labor supply $h_{t}^{i}\in[0,h_{\max}]$: $A_{t}^{h,i}=\{p_{t}^{i},h_{t}^{i}\}.$
@@ 4.2 How effective is LAMP?
LAMP also outperforms language-integrated baselines, demonstrating the advantage of combining MARL with language-guided policy optimization. In S1, using the same backbone and prompt budget, LAMP surpasses the strongest language baseline (ReAct) with +14.8% higher welfare and +14.5% higher reward, while reducing consumption and labor. Under distribution shifts, the advantage remains: in S2 and S3, welfare gains are +1.0% and +10.4%, reward gains are +16.0% and +18.1%, with corresponding reductions in consumption and labor. These results confirm that LAMP’s language-guided coordination improves both stability and efficiency, even in stressed economic conditions.
@@
Isolating Language Effects.
Analysis and Insights.
(2) LAMP addresses this by using LLM reasoning and reflection at each step to extract concise, high‑value insights, which are then passed to the MARL component (e.g., MADDPG). These structured signals—hard for pure data‑driven methods to obtain—are readily produced by pretrained LLMs. Representative examples (above) illustrate the LLM’s clear interpretation of economic variables and targeted reasoning that enhance decision‑making. More examples are shown in Appendix A.
@@
4.3 What drives LAMP’s gains?
Speak Module: Strategy Communication & Opponent Modeling.
Experience Pool: Enhancing Stability and Efficiency.
Reasoning Paths: Trend Tracking & Shock Response.

### v2

# Think, Speak, Decide: Language-Augmented Multi-Agent Reinforcement Learning for Economic Decision-Making

@@ 1 Introduction
To address this challenge, we propose LAMP (Language‑Augmented Multi‑Agent Policy Learning), which integrates LLM‑driven reasoning and reflection over both numerical observations and textual signals to support optimal decision‑making. LAMP follows a unified Think–Speak–Decide pipeline: (1) Think: Agents receive environment observations and generate both short‑term shock analysis and long‑term trend reasoning via an LLM. High‑reward reasoning trajectories are stored in an experience pool for retrieval in similar contexts. The long‑term reasoning is also passed to the Speak module to inform message generation. (2) Speak: Guided by the Think module, each agent formulates multiple candidate public messages. A lightweight attention‑based scorer selects one for broadcast. Other agents parse the message via the LLM, updating their beliefs, trust, and reflective states. These updated reflections are then passed to the Decide module. (3) Decide: The policy network integrates numerical observations, Think’s reasoning outputs, and Speak’s reflections into the RL policy. Under centralized training with a shared critic, agents learn strategies capable of processing reasoning and reflection signals to produce robust, language‑aware economic decisions. We evaluate LAMP in TaxAI and show that it outperforms MARL and LLM-only baselines in both returns and shock robustness. Its reasoning traces explain language-guided choices, aiding insight and policy.
@@
5:   Generate news: $\mathcal{R}_{t}\leftarrow\textbf{Think}(\cdot)_{type}$
@@
11:     $\mathcal{L}_{\text{reason}}(\mathcal{R}_{t},O_{t}^{h,i},\mathcal{H}_{k,t}^{i})$
@@
13:      Generate statement: $v_{t}^{i}\leftarrow\textbf{Speak}(O_{t}^{h,i},\mathcal{R}_{t})$
@@ 3.1 Problem Formulation
We then model the economic decision‑making problem as a partially observable Markov game $\mathcal{M}=\bigl\langle N,\,S,\,O,\,A,\,P,\,R,\,\delta\bigr\rangle,$ where $N=\{g,1,\dots,N_{h}\}$, $\delta\in[0,1)$, and $P$ is the transition kernel induced by $A=A^{g}\times A^{h,1}\times\cdots\times A^{h,N_{h}}.$ At each step, the government’s action is $A_{t}^{g}=\{\tau_{t},\ \xi_{t},\ \tau_{a,t},\ \xi_{a,t},\ r_{t}^{G}\},$ where $\tau_{t}$ and $\xi_{t}$ parameterize the marginal income-tax schedule, $\tau_{a,t}$ and $\xi_{a,t}$ analogously parameterize the marginal asset-tax schedule, and $r_{t}^{G}$ denotes the expenditure-to-output ratio. Each household $i$ selects a savings rate and labor supply $h_{t}^{i}\in[0,h_{\max}]$: $A_{t}^{h,i}=\{p_{t}^{i},h_{t}^{i}\}.$
@@ 4.2 How effective is LAMP?
LAMP also outperforms language-integrated baselines, demonstrating the advantage of combining MARL with language-guided policy optimization. In S1, using the same backbone and prompt budget, LAMP surpasses the strongest language baseline (ReAct) with +14.8% higher welfare and +14.5% higher reward, while reducing consumption and labor. Under distribution shifts, the advantage remains: in S2 and S3, welfare gains are +1.0% and +10.4%, reward gains are +18.1% and +16.0%, with corresponding reductions in consumption and labor. These results confirm that LAMP’s language-guided coordination improves both stability and efficiency, even in stressed economic conditions.
Isolating Language Effects.
Analysis and Insights.
(2) LAMP addresses this by using LLM reasoning and reflection at each step to extract concise, high‑value insights, which are then passed to the MARL component (e.g., MADDPG). These structured signals—hard for pure data‑driven methods to obtain—are readily produced by pretrained LLMs. Representative examples (below) illustrate the LLM’s clear interpretation of economic variables and targeted reasoning that enhance decision‑making. More examples are shown in Appendix A.
4.3 What drives LAMP’s gains?
Speak Module: Strategy Communication & Opponent Modeling.
@@ 4.3 What drives LAMP’s gains?
Experience Pool: Enhancing Stability and Efficiency.
Reasoning Paths: Trend Tracking & Shock Response.